OpenStudy (anonymous):

\[2a_{n-1}-a _{n-2}=3n\] in recurreance relation .. how?????

OpenStudy (amistre64):

ok; what is the question?

OpenStudy (anonymous):

that equ is equal to 3n .. how it is equal to 3n?? that is my question

OpenStudy (amistre64):

well, lets see if we can algebrate it :)

OpenStudy (amistre64):

-A{n-2} = -2A{n-1} +3n ; *-1 A{n-2} = 2A{n-1} - 3n for starters ... right?

OpenStudy (amistre64):

but perhaps the n-1 should ne out front...

OpenStudy (anonymous):

see this if u could get anything for this question = 2(3(n – 1)) – 3(n – 2) = 3n = an.

OpenStudy (amistre64):

2A{n-1} = A{n-2} + 3n /2 A{n-1} = (1/2) A{n-2} + 3n/2

OpenStudy (amistre64):

if we add 1 to all the nth terms we get: A{n} = (1/2) A{n-1} +3(n+1)/2

OpenStudy (amistre64):

and you say the goal is to get to the recursive equation right?

OpenStudy (amistre64):

i dont understand this part: "see this if u could get anything for this question = 2(3(n – 1)) – 3(n – 2) = 3n = an."

OpenStudy (anonymous):

ya actually the left side equation was translated to right side 3n in an recurreance solving ................

OpenStudy (anonymous):

i too actually dont understand anything from that... the question means the equ on left on side since i need to type in equation editor i typed just as question .

OpenStudy (amistre64):

is there a bigger qquestion that this problem comes form? it feels like I am trying to get into it halfway thru.

OpenStudy (anonymous):

no no. not like that

OpenStudy (anonymous):

it is starting only

OpenStudy (anonymous):

wait i will try to google and find anything i can ..

OpenStudy (amistre64):

well, from what you gave me; this is the most sense I can derive from it :)

OpenStudy (amistre64):

A{n} = (1/2) A{n-1} +3(n+1)/2

OpenStudy (anonymous):

wait i will send u the link of that ppt

OpenStudy (amistre64):

ok :)

OpenStudy (anonymous):

this is that ppt see in it

OpenStudy (anonymous):

in slide 4

OpenStudy (amistre64):

i see it; let me read it and see what I can glean :)

OpenStudy (anonymous):

kkkk. waiting for ur reply....

OpenStudy (amistre64):

the first step should be to render a table of values to compare to; right? use the recurence equation to establish the data .... is my first thought

OpenStudy (anonymous):

k..

OpenStudy (amistre64):

if no initial conditions are specified; then a recurrence relation can have multiple solutions; becasue there is no parameters to anchor them to a specific point

OpenStudy (anonymous):

table of values in the sense it must be its initial values??

OpenStudy (amistre64):

if we had initial values; then we could construct a table to match the data to; yes

OpenStudy (amistre64):

suppose I give you an equation like this: y = 2x + d would you be able to graph it?

OpenStudy (anonymous):

ya

OpenStudy (amistre64):

how? since the y intercept is a variable, where are you gonna put it on the graph?

OpenStudy (anonymous):

since it is a recurreance relation they solve it by backward substitution method

OpenStudy (amistre64):

the point they are making in slide 4 is that without initial values; there are many solutions that may exist.

OpenStudy (amistre64):

A{n} = 3n is a possiblility; so is A{n} = 5

OpenStudy (amistre64):

the relation simply states that there is a potential family relations that exhibit the same structure; but differ only by parameters

OpenStudy (amistre64):

just as the line y = 2x + d can have many solutions; it all depends on the value of the y intercept 'd'.

OpenStudy (anonymous):

actually there may be many values. but how that left side = right side eqn . get to this link and see page 2 in that pdf how they have solved that recurreance using backward substitution . http://www.cse.unl.edu/~choueiry/S11-235/files/recitation16.pdf i think we need to do like that to solve left side eqn to bring to right side eqn

OpenStudy (anonymous):

i need to know how they figured out right side eqn from left side eqn??

OpenStudy (anonymous):

by that method can u solve the above eqn on left side and bring it on right side

OpenStudy (amistre64):

well, they didnt really solve it in slide 4; they just determined that #n was a possible solution right? and let me go over it again to make sure im on the same page as them...

OpenStudy (anonymous):

have u seen the pdf ??

OpenStudy (amistre64):

well, yeah, but only if you establish an initial condition; otherwise any A{0} will work

OpenStudy (amistre64):

i saw the pdf, ive even gone thru some of these things before :)

OpenStudy (amistre64):

working backwards you can establish a recurrence equation to determine any nth value and so on

OpenStudy (anonymous):

oh that right. how to establish in the above question and bring 3n??

OpenStudy (amistre64):

A{n} = 2A{n-1} – A{n-2} ; n = 2, 3, 4, … Is A{n} = 3n a solution? 2A{n-1} – A{n-2} ; insert '3(n)' and solve 2 ( 3(n – 1)) – 3(n – 2) ; for n=2, 3(2) 2(3(1)) - 3(0) = 3(2) 2(3) - 0 = 3(2) 2(3) = 3(2) right?

OpenStudy (amistre64):

when n = 3 does it work? 2 ( 3(n – 1)) – 3(n – 2) ; for n=3, 3(3) 2(3(2)) - 3(1) = 3(3) 2(6) - 3 = 3(3) 12 - 3 = 3(3) right?

OpenStudy (amistre64):

Is 3n a solution to the relation? It can be, and it is valid

OpenStudy (anonymous):

ya thanks.............

OpenStudy (amistre64):

We can replace A{n} by 3n and have a valid solution

OpenStudy (amistre64):

slide 3 states what I said ;)

OpenStudy (anonymous):

since we know 3n is the answer we could substitute and prove?? if we have not known how would we have finded the 3n?? that is my question

OpenStudy (amistre64):

'3n' only works because we have no 'initial conditions' in place.

OpenStudy (amistre64):

slide 5 proves that A{n} = 5 is just as valid a solution

OpenStudy (anonymous):

if we have no initial cnd then we have multiple solution?? right that is what written in slide

OpenStudy (amistre64):

that is correct; without knowing where to start from; its impossible to get back there again

OpenStudy (anonymous):

right. are u having any material for studying recurreance relation from basic????

OpenStudy (amistre64):

just whats in the college library over here; oh..and I found a real good teacher online... videos actually. arsdigita, google video arsdigita

OpenStudy (amistre64):

to make a point, all these curves have the same equation except for a y intercept to anchor them. without knowing that initial condition, which graph is the right one?

OpenStudy (anonymous):

ya . correct got your point..... thanks. i will see arsdigita videos.. some other materials ??

OpenStudy (amistre64):

it chopped it .... paste that whole thing in your address bar for the videos

OpenStudy (anonymous):

ya thanks... recurrence comes in discrete mathematics????

OpenStudy (amistre64):

recurrence is alot of discrete math yes :)